OpenStudy (heart_offire):

Choose the graph which represents the solution to the inequality: 1/2x + 8 ≤ 3

OpenStudy (heart_offire):

@welshfella

OpenStudy (welshfella):

have you got a choice of graphs?

OpenStudy (heart_offire):

i do but i cant post it its one of the choices that when i hover over with the mouse it gets bigger

OpenStudy (welshfella):

well lets first simplify the equation if we subtract 8 from both sides we get 1/2 x ≤ -5 can you draw the graph of y = 1/2 x

OpenStudy (welshfella):

it will be a staright line and will pass through the origin because when x = 0 , y = 0 and it will have a slope of 1/2

OpenStudy (heart_offire):

as in now

OpenStudy (welshfella):

?

OpenStudy (heart_offire):

like graph it now?

OpenStudy (welshfella):

I cant draw it somewhow

OpenStudy (heart_offire):

ohhhhh. this isnt a line graph like x,y its a number line

OpenStudy (heart_offire):

my options are number lines

OpenStudy (heart_offire):

sorry for the confusion

OpenStudy (welshfella):

- the area including and below y = -5

OpenStudy (heart_offire):

my options are number lines not graphs

OpenStudy (heart_offire):

when it says graph it means graph on a number line @welshfella

OpenStudy (heart_offire):

@Janu16

OpenStudy (welshfella):

oh ok

OpenStudy (welshfella):

lets solve the inequality 1/2 x + 8 <= 3 subtract 8 from both sides 1/2 x = -5 x = -5 * 2 = -10

OpenStudy (welshfella):

the line graph should look like this